Propanoic acid, CH3CH2COOH, has a pKa of 4.89 at 25 °C. Which of the following is the correct Ka?
1
Expert's answer
2018-05-14T08:17:06-0400
pKa = -lgKa

Ka = 10-pKa

Ka = 10-4.89 = 1.29·10-5